some students simplify 2x^2-32/x-4 as 2x-8/x-4 =2x-8. choose values of x to mshow this is not true. Then simplify it correctly.
Answer by ptaylor(1894)

You need to use parens to define your problem. I could just as easily read your problem as: 2x^2 - (32/x) - 4
Show that 2x-8 is not correct:
Let x=1, then
(2x^2-32)/(x-4)=(2-32)/(1-4)=-30/-3=10;
2x-8=2*1-8=-6
-6 does not equal 10
Simplify correctly:
(2x^2-32)/(x-4)=2(x^2-16)/(x-4);
(x^2-16)=(x+4)(x-4), so
(2x^2-32)/(x-4)=(2(x+4)(x-4))/(x-4); the (x-4)'s cancel
and we have:
2(x+4)=2x+8
does this help????---ptaylor